Desian the Column beamm acior d to Stel conctrultrion maidsa 20 surfaces. weiakt sopst 26 7 Coluwn 斤メed Fixed
Answer #1

area of floor supported by column = 20*20=400 ft2

weight=50 psf

Factored load on column =1.2* (50*400)/1000=24 kips

Column is fixed free

Effective length of column = 2*20=40 feet

Lightest steel section that has compression capacity greater than 24 kips for effectiive length of 40 feet is W8x40